# Tadashi: Enabling AI-Based Automated Code Generation With Guaranteed Correctness

### v1

@@ Acknowledgements.
- Mankowitz et al. (2023) Daniel J. Mankowitz, Andrea Michi, Anton Zhernov, Marco Gelmi, Marco Selvi, Cosmin Paduraru, Edouard Leurent, Shariq Iqbal, Jean-Baptiste Lespiau, Alex Ahern, Thomas Köppe, Kevin Millikin, Stephen Gaffney, Sophie Elster, Jackson Broshear, Chris Gamble, Kieran Milan, Robert Tung, Minjae Hwang, Taylan Cemgil, Mohammadamin Barekatain, Yujia Li, Amol Mandhane, Thomas Hubert, Julian Schrittwieser, Demis Hassabis, Pushmeet Kohli, Martin Riedmiller, Oriol Vinyals, and David Silver. 2023. Faster Sorting Algorithms Discovered Using Deep Reinforcement Learning. Nature 618, 7964 (June 2023), 257–263. https://doi.org/10.1038/s41586-023-06004-9
- Moses et al. (2021) William S Moses, Ruizhe Zhao, Lorenzo Chelini, and Oleksandr Zinenko. 2021. Polygeist: Affine C in MLIR. (2021), 12.
- Munoz et al. (2024) Michael Munoz, Juan Torre, Stefany Lopez, Stiven Herrera, and Christian Uribe. 2024. Comparative Study of AI Code Generation Tools: Quality Assessment and Performance Analysis. LatIA 2 (07 2024), 104. https://doi.org/10.62486/latia2024104
- Mustafa (2022) Dheya Mustafa. 2022. A Survey of Performance Tuning Techniques and Tools for Parallel Applications. IEEE Access 10 (2022), 15036–15055. https://doi.org/10.1109/ACCESS.2022.3147846
@@
- Pouchet et al. (2011) Louis-Noël Pouchet, Uday Bondhugula, Cédric Bastoul, Albert Cohen, J. Ramanujam, P. Sadayappan, and Nicolas Vasilache. 2011. Loop transformations: convexity, pruning and optimization. In Proceedings of the 38th Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages (Austin, Texas, USA) (POPL ’11). Association for Computing Machinery, New York, NY, USA, 549–562. https://doi.org/10.1145/1926385.1926449
@@
- Shido et al. (2019) Yusuke Shido, Yasuaki Kobayashi, Akihiro Yamamoto, Atsushi Miyamoto, and Tadayuki Matsumura. 2019. Automatic Source Code Summarization with Extended Tree-LSTM. In 2019 International Joint Conference on Neural Networks (IJCNN). 1–8. https://doi.org/10.1109/IJCNN.2019.8851751
- Silver et al. (2016) David Silver, Aja Huang, Chris J. Maddison, Arthur Guez, Laurent Sifre, George van den Driessche, Julian Schrittwieser, Ioannis Antonoglou, Veda Panneershelvam, Marc Lanctot, Sander Dieleman, Dominik Grewe, John Nham, Nal Kalchbrenner, Ilya Sutskever, Timothy Lillicrap, Madeleine Leach, Koray Kavukcuoglu, Thore Graepel, and Demis Hassabis. 2016. Mastering the game of Go with deep neural networks and tree search. Nature 529, 7587 (01 Jan 2016), 484–489. https://doi.org/10.1038/nature16961
- Stein et al. (2020) Aviel J. Stein, Levi Kapllani, Spiros Mancoridis, and Rachel Greenstadt. 2020. Exploring Paraphrasing Techniques on Formal Language for Generating Semantics Preserving Source Code Transformations. In 2020 IEEE 14th International Conference on Semantic Computing (ICSC). 242–248. https://doi.org/10.1109/ICSC.2020.00051
@@
- Trümper et al. (2023) Lukas Trümper, Tal Ben-Nun, Philipp Schaad, Alexandru Calotoiu, and Torsten Hoefler. 2023. Performance Embeddings: A Similarity-Based Transfer Tuning Approach to Performance Optimization. In Proceedings of the 37th ACM International Conference on Supercomputing (Orlando, FL, USA) (ICS ’23). Association for Computing Machinery, New York, NY, USA, 50–62. https://doi.org/10.1145/3577193.3593714
- Vasilache et al. (2018) Nicolas Vasilache, Oleksandr Zinenko, Theodoros Theodoridis, Priya Goyal, Zachary DeVito, William S. Moses, Sven Verdoolaege, Andrew Adams, and Albert Cohen. 2018. Tensor Comprehensions: Framework-Agnostic High-Performance Machine Learning Abstractions. CoRR abs/1802.04730 (2018). arXiv:1802.04730 http://arxiv.org/abs/1802.04730
- Verdoolaege (2010) Sven Verdoolaege. 2010. Isl: An Integer Set Library for the Polyhedral Model. In Mathematical Software – ICMS 2010, Komei Fukuda, Joris van der Hoeven, Michael Joswig, and Nobuki Takayama (Eds.). Springer, Berlin, Heidelberg, 299–302. https://doi.org/10.1007/978-3-642-15582-6_49
- Verdoolaege et al. (2013) Sven Verdoolaege, Juan Carlos Juega, Albert Cohen, José Ignacio Gómez, Christian Tenllado, and Francky Catthoor. 2013. Polyhedral Parallel Code Generation for CUDA. ACM Trans. Archit. Code Optim. 9, 4 (Jan. 2013), 54:1–54:23. https://doi.org/10.1145/2400682.2400713
@@
- Verdoolaege et al. ([n. d.]) Sven Verdoolaege, Polly Labs, and KU Leuven. [n. d.]. Presburger Formulas and Polyhedral Compilation. ([n. d.]), 174.
- Wang et al. (2024) Ruotong Wang, Ruijia Cheng, Denae Ford, and Thomas Zimmermann. 2024. Investigating and Designing for Trust in AI-powered Code Generation Tools. In Proceedings of the 2024 ACM Conference on Fairness, Accountability, and Transparency (Rio de Janeiro, Brazil) (FAccT ’24). Association for Computing Machinery, New York, NY, USA, 1475–1493. https://doi.org/10.1145/3630106.3658984
@@
- While and Hingston (2013) Lyndon While and Philip Hingston. 2013. Usefulness of infeasible solutions in evolutionary search: An empirical and mathematical study. In 2013 IEEE Congress on Evolutionary Computation. 1363–1370. https://doi.org/10.1109/CEC.2013.6557723

### v2

@@ Acknowledgements.
- Mankowitz et al. (2023) Daniel J. Mankowitz, Andrea Michi, Anton Zhernov, Marco Gelmi, Marco Selvi, Cosmin Paduraru, Edouard Leurent, Shariq Iqbal, Jean-Baptiste Lespiau, Alex Ahern, Thomas Köppe, Kevin Millikin, Stephen Gaffney, Sophie Elster, Jackson Broshear, Chris Gamble, Kieran Milan, Robert Tung, Minjae Hwang, Taylan Cemgil, Mohammadamin Barekatain, Yujia Li, Amol Mandhane, Thomas Hubert, Julian Schrittwieser, Demis Hassabis, Pushmeet Kohli, Martin Riedmiller, Oriol Vinyals, and David Silver. 2023. Faster Sorting Algorithms Discovered Using Deep Reinforcement Learning. Nature 618, 7964 (June 2023), 257–263. doi:10.1038/s41586-023-06004-9
- Munoz et al. (2024) Michael Munoz, Juan Torre, Stefany Lopez, Stiven Herrera, and Christian Uribe. 2024. Comparative Study of AI Code Generation Tools: Quality Assessment and Performance Analysis. LatIA 2 (07 2024), 104. doi:10.62486/latia2024104
- Mustafa (2022) Dheya Mustafa. 2022. A Survey of Performance Tuning Techniques and Tools for Parallel Applications. IEEE Access 10 (2022), 15036–15055. doi:10.1109/ACCESS.2022.3147846
- OpenAI ([n. d.]) OpenAI. [n. d.]. ChatGPT. https://chat.openai.com
- Ouyang et al. (2024) Anne Ouyang, Simon Guo, and Azalia Mirhoseini. 2024. KernelBench: Can LLMs Write GPU Kernels? In None. https://scalingintelligence.stanford.edu/blogs/kernelbench/
@@
- Pouchet et al. (2011) Louis-Noël Pouchet, Uday Bondhugula, Cédric Bastoul, Albert Cohen, J. Ramanujam, P. Sadayappan, and Nicolas Vasilache. 2011. Loop transformations: convexity, pruning and optimization. In Proceedings of the 38th Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages (Austin, Texas, USA) (POPL ’11). Association for Computing Machinery, New York, NY, USA, 549–562. doi:10.1145/1926385.1926449
@@
- Rogowski et al. (2023) Marcin Rogowski, Samar Aseeri, David Keyes, and Lisandro Dalcin. 2023. Mpi4py.Futures: MPI-Based Asynchronous Task Execution for Python. IEEE Transactions on Parallel and Distributed Systems 34, 2 (Feb. 2023), 611–622. doi:10.1109/TPDS.2022.3225481
@@
- Shido et al. (2019) Yusuke Shido, Yasuaki Kobayashi, Akihiro Yamamoto, Atsushi Miyamoto, and Tadayuki Matsumura. 2019. Automatic Source Code Summarization with Extended Tree-LSTM. In 2019 International Joint Conference on Neural Networks (IJCNN). 1–8. doi:10.1109/IJCNN.2019.8851751
- Silver et al. (2016) David Silver, Aja Huang, Chris J. Maddison, Arthur Guez, Laurent Sifre, George van den Driessche, Julian Schrittwieser, Ioannis Antonoglou, Veda Panneershelvam, Marc Lanctot, Sander Dieleman, Dominik Grewe, John Nham, Nal Kalchbrenner, Ilya Sutskever, Timothy Lillicrap, Madeleine Leach, Koray Kavukcuoglu, Thore Graepel, and Demis Hassabis. 2016. Mastering the game of Go with deep neural networks and tree search. Nature 529, 7587 (01 Jan 2016), 484–489. doi:10.1038/nature16961
- Stein et al. (2020) Aviel J. Stein, Levi Kapllani, Spiros Mancoridis, and Rachel Greenstadt. 2020. Exploring Paraphrasing Techniques on Formal Language for Generating Semantics Preserving Source Code Transformations. In 2020 IEEE 14th International Conference on Semantic Computing (ICSC). 242–248. doi:10.1109/ICSC.2020.00051
- Sullivan (2013) Dennis Michael Sullivan. 2013. Electromagnetic Simulation Using the FDTD Method (second edition ed.). IEEE Press : John Wiley & Sons, Inc, Hoboken, New Jersey.
@@
- Trümper et al. (2023) Lukas Trümper, Tal Ben-Nun, Philipp Schaad, Alexandru Calotoiu, and Torsten Hoefler. 2023. Performance Embeddings: A Similarity-Based Transfer Tuning Approach to Performance Optimization. In Proceedings of the 37th ACM International Conference on Supercomputing (Orlando, FL, USA) (ICS ’23). Association for Computing Machinery, New York, NY, USA, 50–62. doi:10.1145/3577193.3593714
- Vanneschi and Silva (2023) Leonardo Vanneschi and Sara Silva. 2023. Genetic Programming. Springer International Publishing, Cham, 205–257. doi:10.1007/978-3-031-17922-8_8
- Vaughan and Barrett (2015) Courtenay T. Vaughan and Richard F. Barrett. 2015. Enabling Tractable Exploration of the Performance of Adaptive Mesh Refinement. In 2015 IEEE International Conference on Cluster Computing. 746–752. doi:10.1109/CLUSTER.2015.129
- Verdoolaege (2010) Sven Verdoolaege. 2010. Isl: An Integer Set Library for the Polyhedral Model. In Mathematical Software – ICMS 2010, Komei Fukuda, Joris van der Hoeven, Michael Joswig, and Nobuki Takayama (Eds.). Springer, Berlin, Heidelberg, 299–302. doi:10.1007/978-3-642-15582-6_49
- Verdoolaege et al. (2013) Sven Verdoolaege, Juan Carlos Juega, Albert Cohen, José Ignacio Gómez, Christian Tenllado, and Francky Catthoor. 2013. Polyhedral Parallel Code Generation for CUDA. ACM Trans. Archit. Code Optim. 9, 4 (Jan. 2013), 54:1–54:23. doi:10.1145/2400682.2400713
@@
- Verdoolaege et al. (2021) Sven Verdoolaege, Polly Labs, and KU Leuven. 2021. Presburger Formulas and Polyhedral Compilation. (2021), 174.
- Wang et al. (2024) Ruotong Wang, Ruijia Cheng, Denae Ford, and Thomas Zimmermann. 2024. Investigating and Designing for Trust in AI-powered Code Generation Tools. In Proceedings of the 2024 ACM Conference on Fairness, Accountability, and Transparency (Rio de Janeiro, Brazil) (FAccT ’24). Association for Computing Machinery, New York, NY, USA, 1475–1493. doi:10.1145/3630106.3658984
@@
- While and Hingston (2013) Lyndon While and Philip Hingston. 2013. Usefulness of infeasible solutions in evolutionary search: An empirical and mathematical study. In 2013 IEEE Congress on Evolutionary Computation. 1363–1370. doi:10.1109/CEC.2013.6557723